# Learning Beyond Experience: Generalizing to Unseen State Space with Reservoir Computing

### v1

@@ IV.1 Duffing System
To further explore the ability of RCs to generalize, we investigate the effects of training data diversity, e.g., in terms of the range of initial conditions sampled. In Fig. 3d-f, we reduce the half-width, ${\Delta_{0}^{train}}$, of the range from which we select training initial conditions, while holding fixed the half-width of the test initial condition range, ${\Delta_{0}^{test}=10}$. Here, we see that for the RC to reliably predict system behavior in both basins, the initial conditions of the training trajectories must be sufficiently far from the fixed point attractors. For ${\Delta_{0}^{train}=8}$ (Fig. 3d) and ${\Delta_{0}^{train}=6}$ (Fig. 3e), the RC is still able to infer the existence of the unseen attractor and accurately reconstruct a large part of the unseen basin. However, for initial conditions in the outer regions of the test grid, it predicts that the corresponding trajectories converge either to the incorrect attractor (yellow) or to a spurious attractor that is inconsistent with the Duffing system’s true dynamics (white). When the training range is reduced to ${\Delta_{0}^{train}=4}$ (Fig. 3f), the RC fails to identify the unseen fixed point attractor $\bm{A}^{+}\approx(3.16,0)$. Trajectories starting in the white region of Fig. 3d instead converge to a spurious attractor near $x\approx 1.11$ (not shown).In Fig. 7, we present a similar result with RCs trained on fully-observed states of the Duffing system and a similar spurious attractor is clearly visualized.
@@
In Fig. 4, we compare the performance of RCs trained on trajectories from only one basin of attraction, ${B(\bm{A}^{-})}$, to that of RCs trained on trajectories from both basins. To quantify performance, we measure the fraction of trajectories for which the RC predicts the correct attractor,
@@
At every grid point of the heatmaps in Fig. 4, we plot the mean fraction correct, $f_{c}$, averaged over ten independent random draws of the RC’s internal connections, $W_{in}$ and $W_{r}$, and of the initial conditions of the training trajectories.
@@
In Fig. 4a and b, we vary the half-ranges, ${\Delta_{0}^{train}}$ and ${\Delta_{0}^{test}}$, of both the training and test initial conditions. We note that that when ${\Delta_{0}^{train}}$ is sufficiently large ($\Delta_{0}^{train}\gtrsim 7$) and the training time series sample the transient dynamics of the Duffing system far from its attractors, RCs trained on trajectories from only one basin of attraction capture the system’s basin structure just as reliably as those trained on trajectories from both basins. The degradation in performance above the diagonal in both cases, however, demonstrates that RCs have difficulty extrapolating to regions of state space far from their training data – even as those trained on trajectories from only one basin reliably generalize to the unexplored basin. Finally, the white region to the left of Fig. 4a, where the fraction correct is no better than chance, highlights again that RCs trained on only one basin fail to generalize to the unexplored basin if their training signals do not sufficiently sample the transient dynamics of the Duffing system far from its attractors, as illustrated in Fig. 3. In contrast, RCs trained on data from both basins still perform well when $\Delta_{0}^{train}$ is small, so long as $\Delta_{0}^{test}$ is also small.
@@
In Fig. 4c and d, we vary the number of training trajectories, $N_{train}$, while holding the range of test initial conditions fixed at a value large enough that the RC must capture the Duffing system’s transient dynamics well to offer good basin prediction (${\Delta_{0}^{test}=10}$). Here, we see that there is little or no generalization gap. That is, there is no substantial difference in accuracy between RCs trained on data from both basins and those trained on data from only one. In both cases, basin prediction is challenging if the training initial conditions are restricted to a narrow range, even if a large number of training trajectories are available. On the other hand, if the training trajectories are drawn from a wide range of initial conditions and sample well the Duffing system’s transient dynamics, RCs can offer useful basin predictions with very few training trajectories.
IV.2 Magnetic Pendulum
@@
Figure 5: An RC trained on data from only one basin of attraction can capture the magnetic pendulum’s complex basin structure. (a) Basins of attraction of the magnetic pendulum in the plane ${\dot{x}=\dot{y}=0}$. Black crosses mark the positions of the magnets. (b) RC-predicted basin structure using test signals of length ${N_{test}=100}$ data points with initial conditions on a ${300\times 300}$ grid. The training and test signals are partially observed ($\dot{x}$ and $\dot{y}$ withheld); the RC learns to predict only the position components of the pendulum state. All $N_{train}=100$ training trajectories (black dots) are from the pink basin of attraction. Only $f_{spurious}=10^{-4}$ of the predicted trajectories converge to spurious attractors. (c) RC-predicted basins, with all initial conditions for which the RC predicts that the system converges to an incorrect fixed point (real or spurious) colored white.
@@
We now demonstrate that RCs can achieve similar out-of-domain generalization in a multistable system with a more complex basin structure than that of the Duffing system. The magnetic pendulum systemMotter et al. (2013) consists of an iron bob suspended at the end of a pendulum above a plane that contains three magnetic point charges. The magnets sit at the vertices of an equilateral triangle and, when the pendulum hangs straight down, the bob is a height $d$ above the center of this triangle. We choose our coordinate system such that the origin is the triangle’s center and the magnets’ positions are ${(\tilde{x}_{1},\tilde{y}_{1})=(\frac{1}{\sqrt{3}},0)}$, ${(\tilde{x}_{2},\tilde{y}_{2})=(-\frac{1}{2\sqrt{3}},-\frac{1}{2})}$, and ${(\tilde{x}_{3},\tilde{y}_{3})=(-\frac{1}{2\sqrt{3}},\frac{1}{2})}$. Taking the pendulum to be much longer than the distance between magnets (which is one), so that small angle approximations are applicable, the equations of motion are:
@@
is the distance from the bob to the $i^{\text{th}}$ magnet. The pendulum has three stable fixed points, each corresponding to it hanging directly above one of the three magnets. We choose the frequency ${\omega_{0}=0.5}$, damping ${\gamma=0.2}$, and pendulum height ${d=0.2}$, so that the system is dissipative and all trajectories, except for those on the stable manifold of an unstable fixed point at the origin (a set of measure zero), converge to one of the three stable fixed points.
@@
The basin structure of the magnetic pendulum, which we plot in the plane ${\dot{x}=\dot{y}=0}$ in Fig. 5a, is considerably more complex than that of the Duffing system. In fact, while not a true fractal, the basin boundary forms a fractal-like structureMotter et al. (2013) – a so-called ‘slim fractal.’Chen, Nishikawa, and Motter (2017) The resulting sensitivity of the pendulum’s final state to small perturbations makes basin prediction considerably more difficult. Despite the challenge posed by transient chaos, however, we demonstrate in Fig. 5b that an RC trained on trajectories from only one of the magnetic pendulum’s basins of attraction can nonetheless generalize to the other unexplored basins.
Figure 6: RCs may learn some attractors more accurately than others, even as they generalize to unseen basins. (a) to (c) We measure the prediction error (Euclidean distance between the true and predicted trajectories) over time for sample predictions from short test signals (${N_{test}=100}$ data points) in the seen (pink) basin (a) and in the unseen blue (b) and yellow (c) basins. (d) Distributions of the maximum errors over the final $25$ predicted points, calculated over $900$ predictions from test signals with randomly-distributed initial conditions ($310$ pink, $304$ blue, $286$ yellow). We color lines (a to c) according to their predicted basins and histograms (d) according to their true basins. In all panels, dashed lines indicate the threshold distance used to establish convergence, $\varepsilon_{c}=0.25$. We use the same RC and $N_{train}=100$ training trajectories from the pink basin of attraction as in Fig. 5.
@@
The complexity of the pendulum’s basin structure means that small forecast errors can easily push the closed-loop RC system into an incorrect basin of attraction. As a result, we find that to achieve good performance, compared with the Duffing system, our setup for the magnetic pendulum requires: (1) more training trajectories and a more powerful (i.e., larger) reservoir to improve the output-layer accuracy and (2) longer test signals to improve the RC’s initial synchronization. In Fig. 5, for example, we train a reservoir of size ${N_{r}=2500}$ nodes on ${N_{train}=100}$ trajectories, and make predictions using test signals of length ${N_{test}=100}$ data points. (We demonstrate in Fig. 8 that these test signals are still short enough such that predicting the correct basin from the end of the test signal is nearly as difficult as predicting it from the initial condition (start of the test signal). We also show how the RC’s performance varies with the length of the test signals.) Remarkably, despite the fractal-like basin boundaries, the RC is able to provide good predictions not only for the training (pink) basin but also for the two other unseen basins. Still, because of the difficulty of the problem, we do not expect that any RC – even one trained on trajectories from all basins – can reliably predict the correct attractor for initial conditions far from the fixed points. (Even a next-generation reservoir computer with a very strong structural prior requires training data with a high sampling rate to reliably predict the basins of the magnetic pendulum.Zhang and Cornelius (2023)) Fig. 5 illustrates this inherent challenge. The RC-predicted basin structure (Fig. 5b) matches the true basin structure (Fig. 5a) well qualitatively, even as the RC struggles to reliably predict the basins of individual test signals whose initial conditions are in the outer regions of the test grid, where the basin structure is most complex (Fig. 5c). The fraction of test signals for which the RC predicts the correct attractor, $f_{c}\approx 0.67$, however, is still substantially higher than the $47\%$ accuracy achieved by a baseline approach that simply guesses that the pendulum will converge to the magnet nearest to it at the end of the test signal (Fig. 8).
@@
Assessing the qualitative accuracy of the predicted basins in Fig. 5b is analogous to evaluating traditional climate replication in monostable dynamical systems, which focuses on capturing statistical properties of the system over time. Here, our goal is for the statistics of the predicted system behavior – collected over different test signals – to accurately reflect the system’s multistability. Indeed, we see that the RC-predicted trajectories rarely converge to spurious attractors ($f_{spurious}=10^{-4}$, i.e., $9$ of the $90000$ sample predictions in Fig. 5) and the RC broadly captures where extended regions of state space belong to the same basin of attraction and where the basins are more intertwined. (We also illustrate in Fig. 9 that in many scenarios an RC trained on trajectories from only one of the magnetic pendulum’s basins captures the overall basin structure as accurately as an RC trained on data from all three of its basins, similar to what we observe in the Duffing system in Fig. 4.)
In Fig. 6a-c, we plot how the distance between the RC-predicted and true trajectories evolves over time for sample forecasts from each basin of attraction. Predictions from test signals that belong to the seen basin are shown in panel (a) and those from the unseen basins are shown in panels (b) and (c). The color of each line indicates its RC-predicted basin. Interestingly, the RC predicts that trajectories in the yellow basin of attraction converge to a small limit cycle rather than to a fixed point attractor, even as it infers the approximate location of this attractor (within $\varepsilon_{c}=0.25$), correctly identifies the other (blue and pink) fixed-point attractors, and captures the overall structure of all three basins.
Fig. 6d elucidates more thoroughly how the prediction error varies across the attractors. Here, we calculate the maximum distance, $\varepsilon_{25}^{max}$, between the true and predicted pendulum position over the final $25$ predicted time steps (i.e., ${1975\Delta t\leq t\leq 2000\Delta t}$) for each of the $900$ predictions from test signals with initial conditions randomly distributed across ${-1.5\leq x_{0}\leq 1.5}$ and ${-1.5\leq y_{0}\leq 1.5}$. Then, we plot the distributions of ${\varepsilon_{25}^{max}}$ in each basin. Consistent with the limit cycle we observe in the sample predictions (a) to (c), the error of the correct predictions in the yellow basin is the highest among the three basins. In addition, we see that the RC learns the near-attractor behavior in the seen pink basin more accurately than in the unseen blue basin.
Overall, Fig. 6 demonstrates that an RC can generalize in an operationally useful way, even without achieving equal accuracy in the seen and unseen domains. Moreover, it suggests that an RC can learn the system dynamics in a manner that allows it to generalize to unseen regions of state space without strictly capturing system symmetries.

### v2

@@ IV.1 Intertwined Basins in the Duffing System
To further explore the ability of RCs to generalize, we investigate the effects of training data diversity, e.g., in terms of the range of initial conditions sampled. In Fig.˜3b-c, we reduce the half-width, $\Delta_{0}^{train}$}, of the range from which we select training initial conditions, while holding fixed the half-width of the test initial condition range, $\Delta_{0}^{test}=10$}. Here, we see that for the RC to reliably predict system behavior in both basins, the initial conditions of the training trajectories must be sufficiently far from the fixed-point attractors. For $\Delta_{0}^{train}=7$} (Fig.˜3b), the RC is still able to infer the existence of the unseen attractor and accurately reconstruct a large part of the unseen basin. However, for some initial conditions in the outer regions of the test grid, it predicts that the corresponding trajectories converge to the incorrect attractor (yellow). When the training range is reduced to $\Delta_{0}^{train}=4$} (Fig.˜3c), the RC fails to identify the unseen fixed-point attractor $\bm{A}^{+}\approx(3.16,0)$. Trajectories starting in the white regions of Fig.˜3c instead converge to spurious fixed-point attractors near $x\approx-8.4$ and $x\approx 1.75$ (not shown) that are inconsistent with the Duffing system’s true dynamics. In the supplementary material (Fig.˜S1), we present a similar result with RCs trained on fully-observed states of the Duffing system and a similar spurious attractor is clearly visualized.
We demonstrate in Fig.˜3f-g, that the RC does not rely on the Duffing system’s rotational symmetry to generalize to the unexplored basin. Specifically, we break the symmetry by adding a constant external forcing, $F_{0}=1$}, to Eq.˜13b. This change also moves the attracting fixed points to $\bm{A}^{-}_{F}\approx\left(-2.42,0\right)$} and $\bm{A}^{+}_{F}\approx\left(3.58,0\right)$}. We then train the RC on $N_{train}=10$} partially-observed trajectories from a single basin of attraction and make predictions from short test signals ($N_{test}=10$}) in both basins of attraction, as before. The RC still recovers the system behavior in both basins, whether we train it on data from the smaller blue basin (panel f) or from the larger pink basin (panel g).
@@
In Fig.˜4, we compare the performance of RCs trained on trajectories from only one basin of attraction of the unforced Duffing system, $B(\bm{A}^{-})$}, to that of RCs trained on trajectories from both basins. To quantify performance, we measure the fraction of trajectories for which the RC predicts the correct attractor,
@@
At every grid point of the heatmaps in Fig.˜4, we plot the mean fraction correct, $f_{c}$, averaged over ten independent random draws of the RC’s internal connections, $W_{in}$ and $W_{r}$, and of the initial conditions of the training trajectories.
@@
In Fig.˜4a and b, we vary the half-ranges, $\Delta_{0}^{train}$} and $\Delta_{0}^{test}$}, of both the training and test initial conditions. We note that that when $\Delta_{0}^{train}$} is sufficiently large ($\Delta_{0}^{train}\gtrsim 7$) and the training time series sample the transient dynamics of the Duffing system far from its attractors, RCs trained on trajectories from only one basin of attraction capture the system’s basin structure just as reliably as those trained on trajectories from both basins. The degradation in performance above the diagonal in both cases, however, demonstrates that RCs have difficulty extrapolating to regions of state space far from their training data – even as those trained on trajectories from only one basin reliably generalize to the unexplored basin. Finally, the white region to the left of Fig.˜4a, where the fraction correct is no better than chance, highlights again that RCs trained on only one basin fail to generalize to the unexplored basin if their training signals do not sufficiently sample the transient dynamics of the Duffing system far from its attractors, as illustrated in Fig.˜3. In contrast, RCs trained on data from both basins still perform well when $\Delta_{0}^{train}$ is small, so long as $\Delta_{0}^{test}$ is also small.
@@
In Fig.˜4c and d, we vary the number of training trajectories, $N_{train}$, while holding the range of test initial conditions fixed at $\Delta_{0}^{test}=10$} (for consistency with Fig.˜3, and because it is large enough that the RC must capture the Duffing system’s transient dynamics well to offer good basin prediction). Here, we see that there is little or no generalization gap. That is, there is no substantial difference in accuracy between RCs trained on data from both basins and those trained on data from only one. In both cases, basin prediction is challenging if the training initial conditions are restricted to a narrow range, even if a large number of training trajectories is available. On the other hand, if the training trajectories are drawn from a wide range of initial conditions and sample well the Duffing system’s transient dynamics, RCs can offer useful basin predictions with very few training trajectories.
We demonstrate in the supplementary material (Fig.˜S2) that RCs can still reliably generalize to an unseen basin of the unforced Duffing system when moderate stochastic forcing (dynamical/process noise) influences the evolution of the training trajectories.
IV.2 Segregated Basins in a Multi-well System
In the Duffing system, the basins of attraction for the two fixed-point attractors are substantially intertwined. When the training initial conditions span a wide range of phase space, some of the resulting trajectories from the observed basin spend time close to regions belonging to the unseen basin of attraction, providing useful information for generalization. In this section, we investigate how well our RC setup can generalize in a system with segregated basins of attraction, where training trajectories may start near basin boundaries but quickly move away from them, making the extrapolation task potentially more challenging.
Figure 5: RC identification of unseen attractors in a multi-well system with segregated basins. We train an RC of $N_{r}=200$} nodes on $N_{train}=25$} fully-observed trajectories (gray lines) from two (a, b) or just one (c) of the multi-well system’s basins of attraction (left). Then, we forecast the system’s evolution from $100$ initial conditions (dots) and corresponding short test signals (thick lines, $N_{test}=5$} observations) in two cases: when the inputs to RC are standardized (center), and when the inputs to the RC are not standardized (right). We color predictions (thin lines) orange if they go to an incorrect attractor, and according to their true basin otherwise.
Specifically, we employ a simple multi-well potential with four segregated basins of attraction. The dynamics are governed by two independent differential equations:
$\dot{x}=\frac{1}{2}x(1-x^{2}),$ (16a)
$\dot{y}=\frac{1}{2}y(1-y^{2}),$ (16b)
which yield four stable fixed-point attractors at the corners of the unit square and an unstable fixed point at the origin.
We demonstrate in Fig.˜5 that our RC setup can generalize to unseen segregated basins in the multi-well system. However, the quality of this generalization is somewhat influenced by whether inputs to the reservoir are standardized based on the training distribution (a common practice in machine learning). Standardization schemes aim to ensure that each component of the inputs has a similar scale, which helps prevent a subset of inputs from dominating the dynamics and promotes more effective use of the system’s computational capacity. However, standardization schemes that are based on the training distribution implicitly bias the model to perform well on data with similar distributions (i.e., prioritize in-distribution performance). This bias may not hinder generalization to unseen intertwined basins of attraction, but we show in Fig.˜5 that it can slightly degrade performance in unseen segregated basins of attraction, where the mean and range of the test data differ substantially from the training data.
In the central column of Fig.˜5, we use the input strength range $\sigma=1.0$ (Table˜3), and we standardize all inputs such that the training inputs have component-wise mean zero and maximum absolute value one, as before. (It is also common to standardize inputs to have unit variance over the training data, but we use the maximum absolute value in this paper so that the standardization is not affected by long trajectories that are near a fixed point.) In the right-hand column, we do not standardize the training or test signal inputs and we scale the input strength range, $\sigma=0.25$, accordingly. When the range of the training trajectories approximately spans the range of the test initial conditions, $4\lesssim x\lesssim 4$} and $4\lesssim y\lesssim 4$} (row a), the RC performs similarly with standardized and raw inputs. When the training trajectories sample only the left two basins of attraction, $x<0$ (row b), the RC again generalizes well in both cases, but captures the locations of the unseen fixed points at x=1{x=1} slightly more accurately when the input time series are not standardized. Finally, if the training data sample only one basin of attraction (row c), the RC identifies all of the unseen attractors when the inputs are not standardized (raw) but fails to infer one of the unseen fixed-point attractors when the inputs to the reservoir are standardized.
IV.3 Complex Basin Structure in a Magnetic Pendulum System
@@
Figure 6: An RC trained on data from only one basin can capture the magnetic pendulum’s complex basin structure. (a) Basins of attraction of the magnetic pendulum in the plane $\dot{x}=\dot{y}=0$}. Black crosses mark the positions of the magnets. (b) RC-predicted basin structure using test signals of length $N_{test}=100$} data points with initial conditions on a $300\times 300$} grid. The training and test signals are partially observed ($\dot{x}$ and $\dot{y}$ withheld); the RC learns to predict only the position components of the pendulum state. All $N_{train}=100$ training trajectories (black dots) are from the pink basin of attraction. Only $f_{spurious}\approx 8\times 10^{-4}$} of the predicted trajectories converge to spurious attractors (green). (c) RC-predicted basins, with all initial conditions for which the RC predicts that the system converges to an incorrect fixed point (real or spurious) colored white.
@@
We now demonstrate that RCs can achieve similar out-of-domain generalization in a multistable system with a more complex basin structure than those of the Duffing or multi-well systems. The magnetic pendulum systemMotter et al. (2013) consists of an iron bob suspended at the end of a pendulum above a plane that contains three magnetic point charges. The magnets sit at the vertices of an equilateral triangle and, when the pendulum hangs straight down, the bob is a height $d$ above the center of this triangle. We choose our coordinate system such that the origin is the triangle’s center and the magnets’ positions are $(\tilde{x}_{1},\tilde{y}_{1})=(\frac{1}{\sqrt{3}},0)$}, $(\tilde{x}_{2},\tilde{y}_{2})=(-\frac{1}{2\sqrt{3}},-\frac{1}{2})$}, and $(\tilde{x}_{3},\tilde{y}_{3})=(-\frac{1}{2\sqrt{3}},\frac{1}{2})$}. Taking the pendulum to be much longer than the distance between magnets (which is one), so that small angle approximations are applicable, the equations of motion are:
@@
is the distance from the bob to the $i^{\text{th}}$ magnet. The pendulum has three stable fixed points, each corresponding to it hanging directly above one of the three magnets. We choose the frequency $\omega_{0}=0.5$}, damping $\gamma=0.2$}, and pendulum height d=0.2{d=0.2}, so that the system is dissipative and all trajectories, except for those on the stable manifold of an unstable fixed point at the origin (a set of measure zero), converge to one of the three stable fixed points.
@@
The basin structure of the magnetic pendulum, which we plot in the plane $\dot{x}=\dot{y}=0$} in Fig.˜6a, is considerably more complex than that of the Duffing system. In fact, while not a true fractal, the basin boundary forms a fractal-like structureMotter et al. (2013) – a so-called ‘slim fractal.’Chen, Nishikawa, and Motter (2017) The resulting sensitivity of the pendulum’s final state to small perturbations makes basin prediction considerably more difficult. Despite the challenge posed by transient chaos, however, we demonstrate in Fig.˜6b that an RC trained on trajectories from only one of the magnetic pendulum’s basins of attraction can nonetheless generalize to the other unexplored basins.
@@
The complexity of the pendulum’s basin structure means that small forecast errors can easily push the closed-loop RC system into an incorrect basin of attraction. As a result, we find that to achieve good performance, compared with the Duffing and multi-well systems, our setup for the magnetic pendulum requires: (1) more training trajectories and a more powerful (i.e., larger) reservoir to improve the output-layer accuracy and (2) longer test signals to improve the RC’s initial synchronization. In Fig.˜6, for example, we train a reservoir of size $N_{r}=2500$} nodes on $N_{train}=100$} trajectories, and make predictions using test signals of length $N_{test}=100$} data points. (We demonstrate in Fig.˜S3 of the supplementary material that these test signals are still short enough such that predicting the correct basin from the end of the test signal is nearly as difficult as predicting it from the initial condition (start of the test signal). We also show how the RC’s performance varies with the length of the test signals.) Remarkably, despite the fractal-like basin boundaries, the RC is able to provide good predictions not only for the training (pink) basin but also for the two other unseen basins. Still, because of the difficulty of the problem, we do not expect that any RC – even one trained on trajectories from all basins – can reliably predict the correct attractor for initial conditions far from the fixed points. (Even a next-generation reservoir computer with a very strong structural prior requires training data with a high sampling rate to reliably predict the basins of the magnetic pendulum.Zhang and Cornelius (2023)) Fig.˜6 illustrates this inherent challenge. The RC-predicted basin structure (Fig.˜6b) matches the true basin structure (Fig.˜6a) well qualitatively, even as the RC struggles to reliably predict the basins of individual test signals whose initial conditions are in the outer regions of the test grid, where the basin structure is most complex (Fig.˜6c). The fraction of test signals for which the RC predicts the correct attractor, $f_{c}\approx 0.66$, however, is still substantially higher than the baseline fraction of $0.47$ achieved by an approach that simply guesses that the pendulum will converge to the magnet nearest to it at the end of the test signal (supplementary material Fig.˜S3).
@@
Assessing the qualitative accuracy of the predicted basins in Fig.˜6b is analogous to evaluating traditional climate replication in monostable dynamical systems, which focuses on capturing statistical properties of the system over time. Here, our goal is for the statistics of the predicted system behavior – collected over different test signals – to accurately reflect the system’s multistability. Indeed, we see that the RC-predicted trajectories rarely converge to spurious attractors ($f_{spurious}\approx 8\times 10^{-4}$, i.e., $75$ of the $90000$ sample predictions in Fig.˜6) and the RC broadly captures where extended regions of state space belong to the same basin of attraction and where the basins are more intertwined.
Table 1: Per-basin RC performance in predicting the correct attractor in Fig.˜6. Baseline: a predictor that always guesses the magnet nearest the pendulum at the end of the test signal.
Basin Fraction Correct False Neg. Rate False Pos. Rate
Pink 0.66 0.34 0.17
Blue 0.65 0.35 0.17
Yellow 0.66 0.34 0.18
Baseline 0.47
Figure 7: RC reconstruction of an unseen chaotic attractor in a multistable Lorenz-like system. We train an RC with $N_{r}=500$} nodes on a single fully-observed trajectory (dashed line) of the Lorenz-like system. (a) We forecast the system’s evolution from $36$ short test signals ($N_{test}=50$} observations, test signal trajectories not shown for clarity), each starting from a random initial condition. Forecasts start from the endpoints of these test signals, and each predicted trajectory is colored according to the true basin of its initial condition. (b) We make predictions from shorter test signals ($N_{test}=5$}) with initial conditions forming a $100\times 100$} grid in the $x=0$ plane, and plot the RC-predicted basin structure. Yellow and white, respectively, indicate initial conditions for which the prediction converges to the incorrect attractor and those for which the prediction does not accurately resemble either true attractor ($\Delta_{KL}(P_{A_{i}},P_{\hat{\bm{u}}})>\epsilon_{c}$} for i=1,2{i=1,2}). (c) Distribution of the KL divergence between the predicted and true attractors, computed over all predictions in panel (b). Each histogram is formed from predictions associated with test signals from the same true basin of attraction and is colored accordingly. The vertical dotted line indicates the convergence threshold, $\varepsilon_{c}=1$}, and the vertical dashed line marks the KL divergence between the true attractors, estimated as $\frac{1}{2}\left(\Delta_{KL}(P_{A_{1}},P_{A_{2}})+\Delta_{KL}(P_{A_{2}},P_{A_{1}})\right)$}.
At the per-basin level, the fraction correct ($f_{c}$), false negative rate ($1-f_{c}$), and false positive rate are summarized in Table˜1. These results, coupled with Fig.˜S4 of the supplementary material, indicate that RC performance in predicting basin membership is comparable for seen and unseen basins, and that, in many cases, an RC trained on trajectories from only one basin of the magnetic pendulum still recovers the overall basin structure almost as well as an RC trained on all three basins – consistent with what we find for the Duffing system (Fig.˜4). However, even when the RC is successful at predicting basin membership, it may still miss some of the detailed dynamics of the unseen basins (see supplementary material Fig.˜S5). Overall, these findings suggest that an RC can learn the system dynamics in a way that enables generalization to unseen regions of state space in an operationally useful manner, even without strictly capturing system symmetries.
IV.4 Coexisting Chaotic Attractors in a Lorenz-like System
The Duffing, multi-well, and magnetic pendulum systems are dissipative, with all trajectories converging to fixed-point attractors. In many multistable systems, however, the system state never reaches a fixed point and instead converges to a limit cycle or chaotic attractor, along which it evolves indefinitely. In this section, we highlight a case in which an RC captures the existence of an unseen chaotic attractor.
Following Lü et al.,Lü, Chen, and Cheng (2004) we study a multi-stable Lorenz-like system that is governed by three coupled ordinary differential equations:
$\dot{x}=-\frac{ab}{a+b}x-yz+c,$ (19a)
$\dot{y}=ay+xz,$ (19b)
$\dot{z}=bz+xy.$ (19c)
When a=10{a=-10}, b=4{b=-4}, and c=18.1{c=18.1}, the system has two chaotic attractors. In Fig.˜7a, we train an RC on a single long trajectory (black dashed line), which approaches and then evolves along one of the attractors. Then we show that the RC captures the existence and structures of both the seen attractor (the blue lobe, $A_{1}$}) and the unseen attractor (the pink lobe, $A_{2}$}). To obtain a clear image of the predicted chaotic attractors in Fig.˜7a, we make predictions using test signals of $N_{test}=50$} data points and omit the test signals themselves from the figure. These test signals are, however, long enough that each prediction starts when the system state has already reached one of the chaotic attractors.
In Fig.˜7b, we reduce the length of the test signals to $N_{test}=5$} observations so that, for most initial conditions, the system is still far from the attractors when prediction begins. Then we plot the RC-predicted basin structure for test signals with initial conditions in the plane $x=0$. The RC accurately captures the basin structure of the multistable Lorenz-like system. In fact, the RC predicts the correct attractor for almost all initial conditions in the $100\times 100$} test grid, indicating that it can accurately capture the system’s transient behavior en route to the chaotic attractors (in addition to reconstructing the chaotic attractors themselves).
The convergence criteria used for the Duffing, multi-well, and magnetic pendulum systems in Sections˜IV.1, IV.2 and IV.3 do not apply to the Lorenz-like system because its attractors are not fixed points. Instead, to determine whether a trajectory $\bm{u}(t)$} converges to one of the system’s chaotic attractors, we compute the Kullback-Liebler (KL) divergence of the empirical state distribution, $P_{\bm{u}}(\bm{v})$, over the final 500 time steps of $\bm{u}(t)$} relative to the distribution $P_{A}(\bm{v})$ from a reference trajectory on the attractor:
$\Delta_{KL}\left(P_{A},P_{\bm{u}}\right)=\int_{\bm{v}\,\in\,\mathbb{R}^{3}}P_{A}(\bm{v})\log\frac{P_{A}(\bm{v)}}{P_{\bm{u}}(\bm{v})}d\bm{v}.$ (20)
Following previous work,Zhang and Gilpin (2025); Hess et al. (2023) we use Gaussian mixture models to estimate the distributions $P_{\bm{u}}(\bm{v})$} and $P_{A}(\bm{v})$} from the predicted and reference trajectories. If the KL divergence of the predicted distribution relative to the true distribution on the attractor $A$ is less than a threshold value, $\Delta_{KL}(P_{A},P_{\bm{u}})<\varepsilon_{c}$}, we say that the predicted trajectory converges to the attractor $A$.
In Fig.˜7c, we use the KL divergence to assess the quality of the reconstructed attractors in the seen and unseen basins. The RC achieves comparable accuracy on the unseen (pink) attractor as on the seen (blue) attractor, with the KL divergence of the predicted attractor relative to the true attractor in both cases substantially lower than the KL divergence between the two true attractors (vertical dashed line).